OpenStudy (anonymous):

Sam wants to leave an 18% tip for his dinner. The bill is for $23.50. Which equation could be used to find the total amount that Sam should pay? a,23.5 = 0.18x b,x= (0.18)(23.5) c,x= (23.5)(1.18) d,1.18 = 23.5x please help me:(

OpenStudy (nurali):

(23.5)(1.18)

OpenStudy (nurali):

if correct please click best response

OpenStudy (anonymous):

Sam would be adding 18% of the bill ($23.50) to the bill ($23.50). Let what Sam has to pay be x. Then the equation is x=(.18)*(23.50)+23.50 ==> x=(1+.18)*23.50 ==> x=(1.18)*(23.50) So, the answer is c.

OpenStudy (anonymous):

so wich one is right lol? the other person says its a

OpenStudy (anonymous):

never mind its all good lol